OpenStudy (anonymous):

using the two functions listed below, insert numbers in place of the letters a,b,c and d so that f(x) and g(x) are inverses. (i am going to put the functions in the second comment)

OpenStudy (anonymous):

\[f(x)= \frac{ x+a }{ b }\] \[g(x)=cx-d\]

OpenStudy (anonymous):

@BeautifulMystery. @mathwizzard3 @parker.goodbar @readergirl12

OpenStudy (precal):

Why not just put numbers in f(x), find the inverse and then call it g(x)?

OpenStudy (anonymous):

First, you need to find the inverse of either f(x) or g(x). Do you remember how to do that?